Differential Equations
Need help on understanding how to find General Solutions...

General Solution of Bernoulli equation

y' not y^1

thank you!!
Answer #1

* Gen By + 52 y = 7x3 Jy + 5 = 7 y 3 - let.qt — Replacing 0 and 1 in ② -- 10. lntal 2) t-ot - 7x2 2 2 2) t- 10 x = -14/22 N£nto = 1683 te - x = 168 9² + cy²3 - [multietying soith y 213) General solution is 2²_1689²_cg ²x² ao (ewhere e is an anbitarThus First the Bernoulli's differential equation is converted into linear differential equation with the change of variable and then the linear differential equation is solved.